### v2

# Enhancing Table Recognition with Vision LLMs: A Benchmark and Neighbor-Guided Toolchain Reasoner

Abstract.
Pre-trained foundation models have recently significantly progressed in structured table understanding and reasoning. However, despite advancements in areas such as table semantic understanding and table question answering, recognizing the structure and content of unstructured tables using Vision Large Language Models (VLLMs) remains under-explored. In this work, we address this research gap by employing VLLMs in a training-free reasoning paradigm. First, we design a benchmark with various hierarchical dimensions relevant to table recognition. Subsequently, we conduct in-depth evaluations using pre-trained VLLMs, finding that low-quality image input is a significant bottleneck in the recognition process. Drawing inspiration from these findings, we propose the Neighbor-Guided Toolchain Reasoner (NGTR) framework, which is characterized by integrating multiple lightweight models for low-level visual processing operations aimed at mitigating issues with low-quality input images. Specifically, we utilize a neighbor retrieval mechanism to guide the generation of multiple tool invocation plans, transferring tool selection experiences from similar neighbors to the given input, thereby facilitating suitable tool selection. Additionally, we introduce a reflection module to supervise the tool invocation process. Extensive experiments on public table recognition datasets demonstrate that our approach significantly enhances the recognition capabilities of the vanilla VLLMs. We believe that the designed benchmark and the proposed NGTR framework could provide an alternative solution in table recognition [^1] Our code is open-sourced and available at https://github.com/lqzxt/NGTR.
Vision Large Language Models, Table Recognition, Tool Invocation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Tables are ubiquitous for organizing and communicating structured information across diverse domains, ranging from scientific literature and business reports to web pages and financial documents (zhang2020web, ; zheng2021global, ). They encapsulate a wealth of information essential for numerous applications, such as knowledge discovery, decision support, and data-driven analytics (shwartz2022tabular, ; wang2024tabletime, ). In the context of intelligent table applications, one fundamental yet challenging task is table recognition (zanibbi2004survey, ): converting image-based table representations into structured digital formats. Over the years, substantial efforts (salaheldin2024deep, ) have been made to tackle this problem, introducing various approaches to address challenges such as segmentation techniques and object detection methods.
Figure 1. Comparison of modeling paradigms: domain-specific lightweight models vs. universal pre-trained VLLMs.
Recently, the advent of Large Language Models (LLMs) (chang2024survey, ) and Vision Large Language Models (VLLMs) (yin2024survey, ) has revolutionized natural language processing and computer vision, respectively. For LLMs, their ability to process serialized data has facilitated numerous tabular data mining tasks, such as table-to-text generation (liu2018table, ; cheng2021hitab, ; guo2024adapting, ), table question answering (wang2024chain, ; gao2023text, ; mao2024potable, ), and table semantic understanding (deng2022turl, ; feuer2023archetype, ; trabelsi2020semantic, ). Meanwhile, several VLLM-based methods have emerged to bypass traditional OCR pipelines for visual table analysis and understanding (zheng2024multimodaltableunderstanding, ; hu2024mplug, ).
Despite these advancements, our investigation reveals a noticeable gap: the application of VLLMs to table recognition remains underexplored. This task serves as a foundational building block for table-related applications. Some existing work (chen2023tablevlm, ; zhang2024multi, ) has focused on pre-training or fine-tuning VLLMs to accomplish this task. However, Fine-tuning VLLMs for specific tasks is often computationally expensive and risks catastrophic forgetting of general capabilities. To address this, we focus on leveraging a prompt-based paradigm for table recognition using pre-trained VLLMs, exploring a generative approach without requiring additional fine-tuning. Recognizing the absence of dedicated benchmarks in this domain, we design an evaluation framework based on a hierarchical design philosophy (sui2024table, ; cheng2023formertime, ; liu2024generative, ), comprising recognition tasks for table recognition. Through extensive evaluations, we identify a critical bottleneck: low-quality input images significantly hinder the table recognition capabilities of the evaluated VLLMs. This is understandable given the various characteristics associated with different table styles.
To overcome this limitation, we propose the Neighbor-Guided Toolchain Reasoner (NGTR) framework with the goal of effective table recognition. One of the main characteristics of the NGTR framework is that it integrates the capacity of lightweight models and the strategy of retrieval-augmented generation to enhance input image quality and guide structured data recognition. Specifically, we propose a preprocessing toolkit equipped with various lightweight models designed to enhance input image quality. For each input instance, we retrieve a similar neighbor from the training data, and the experience gained from these neighbors is used to inform the guidance for the generation of tool invocation plans. This experience based guidance helps select and integrate suitable tools for the given input. Furthermore, we incorporate a reflection-driven tool selection module at each step of tool invocation to iteratively refine table recognition outputs. This enables VLLMs to produce higher-quality structured data.
To validate the effectiveness of the proposed NGTR framework, we conduct extensive experiments on multiple public table recognition datasets. The key observations are as follows: (1) Our NGTR framework significantly enhances the table recognition performance of naive VLLM-based approaches; (2) While VLLMs achieve competitive accuracy on specific datasets compared to traditional models, a noticeable performance gap remains in favor of traditional models. Nonetheless, we have preliminarily revealed the performance boundaries of VLLMs in several representative table recognition datasets. As is shown in Figure 1, the VLLM-based table recognition approach demonstrates the capability for universal modeling. This method facilitates a paradigm shift in design objectives from a model-centric to a data-centric focus, presenting significant potential for further exploration. We hope this work will inspire more research efforts in the future.
In summary, the contributions of this paper are as follows:
  • We pioneer the exploration of VLLM-based table recognition tasks by introducing a comprehensive benchmark with hierarchical evaluation settings.
  • We propose the NGTR framework to address critical bottlenecks in table recognition, such as low-quality input images.
  • We conduct extensive experiments to report the promising performance and potential of VLLMs for table recognition, along with interesting observations that highlight areas for future research.
2. Related Work
Table Recognition. Earlier table recognition (TR) methods predominantly rely on heuristic rules (TR-Heu-1, ; TR-Heu-2, ; TR-Heu-3, ) or statistical learning techniques (TR-Stat-1, ; TR-Stat-2, ), which were generally effective on simple and standard structured tables. However, these approaches rely heavily on handcrafted features or implicit rules and show limited generalization ability. In the era of deep learning, numerous studies have made impressive progress in handling more intricate and heterogeneous table structures. Top-down methods (TR-TD-1, ; TR-TD-2, ; TR-TD-3, ; qin2024semv3, ) first predict table lines and boundaries to infer the overall structure, and then locate cells based on row-column intersections. Bottom-up methods (zheng2021global, ; qiao2021lgpma, ; xing2023lore, ) first identify table cells with object detection models (OD-1, ; OD-2, ), and then predict the cell relations to organize the row-column structures to form the overall tables. These methods follow an explicit two-stage learning paradigm with relatively strong transferability and explainability, yet the risks of ambiguous contents and boundless structures may cause unstable and incorrect predictions.
Recently, sequence-based methods (zhong2020image, ; TR-Seq-2, ; TR-Seq-3, ; TR-Seq-4, ) have been widely explored to directly generate markup sequence defining structures with specific decoders. Although these approaches require a massive of training data and computing resources, they have demonstrated substantial potential to unify visual-text parsing tasks (wan2024omniparser, ).
Large Language Models. In recent years, LLMs have made significant progress in the field of natural language processing. In particular, LLMs such as GPT (radford2018improving, ) and Llama (touvron2023llama, ) have demonstrated exceptional performance in tasks such as multi-task learning (chen2024multi, ), zero-shot learning (kojima2022large, ), and text generation (li2024pre, ). These models have not only broken through the limitations of traditional technologies in generating natural language text, but have also shown capabilities in reasoning (wei2022chain, ; cheng2024advancing, ; shinn2024reflexion, ) and planning (guan2023leveraging, ; gou2023critic, ). Meanwhile, VLLMs combine the capabilities of visual and language understanding, enabling LLMs to process visual information. For multimodal understanding scenarios (e.g., scene text recognition (STR, ), visual question answering (VQA, )), VLLMs have been widely validated as effective (Image2LLM, ; UReader, ; TextMonkey, ). With continuous progress in text-rich scenarios, some studies (zheng2024multimodaltableunderstanding, ; zhao2024tabpedia, ; chen2023tablevlm, ) have also focused on enabling VLLMs to handle multimodal table understanding tasks.
Despite their promising success in a wide range of domains, VLLMs applied to table recognition remain under-evaluated and under-explored. Our study presents a comprehensive benchmark for VLLM-based table recognition evaluation. Subsequently, we propose a novel framework to address the bottleneck of VLLMs, thereby enhancing their capabilities in table recognition.
3. Problem Definition and Proposed Benchmark
3.1. Problem Definition
We employ the generation paradigm of VLLMs to address the table recognition (TR) task, which is formulated as a format mapping problem from images to sequences. Formally, given a TR dataset $\mathcal{D}=\{(I^{i},H^{i})\}_{i=1}^{n}$ with $n$ samples, we predict the corresponding structured form $H^{i}$ for each table image $I^{i}$. Specifically, we provide the image table $I^{i}$ along with a prompt $P$ as input to the VLLMs, which generates the structured data form $\hat{H}^{i}=\operatorname{VLLM}(P,I^{i})$.
3.2. Benchmark Evaluation Setup
This section proposes an evaluation benchmark for TR based on VLLMs, outlining the hierarchical recognition tasks to assess their performance and the evaluation setup, including the recognition tasks and VLLMs being evaluated. Details of the evaluation metrics and datasets are discussed in Section 5.1.
Table 1. Descriptions of the proposed hierarchical recognition tasks.
Granularity Recognition Task Description
Table-level
Visual Table Size
Detection
Get the number of rows and columns.
Row-level
Row Index-based
Data Recognition
Get the content list of a specific row.
Column-level
Column Index-based
Data Recognition
Get the content list of a specific column.
Cell-level
Merged Cell
Detection
Get contents of all merged cells.
Content-based
Cell recognition
Get the location of specific cell content.
Index-based
Cell Recognition
Get the cell content of specific location.
3.2.1. Recognition Task Design
We design several hierarchical recognition tasks to conduct a more in-depth assessment of VLLMs’ table recognition capability. In addition to the general table recognition task, we extend our approach by designing recognition tasks that focus on the structural composition of the table, including the cell-level, row-level, column-level, and global table-level. This hierarchical recognition tasks design comprehensively assesses VLLM’s table recognition capability at different levels. Table 1 presents the details of the hierarchical recognition tasks.
Cell-level. We evaluate the cell-level recognition capability of VLLMs within three specific recognition tasks. Merged cell detection task aims to recognize cells that span multiple rows or columns in the table. Formally, given an image $I$, this task aims to recognize $\{c_{1},c_{2},\ldots,c_{n}\}$ that represents the textual content of all merged cells in the table. Content/index-based cell recognition tasks evaluate the structural and content recognition capabilities of VLLMs, which are crucial for assessing whether VLLMs could perform well in fine-grained table recognition. Formally, define an image $I$, the cell content $c$, and the cell’s row and column indices $(R,C)$. The Content-based cell recognition task identifies the logical location $(R,C)$ of the target cell, and the index-based cell recognition task extracts the content of the target cell $c_{(R,C)}$.
Row/Column-level. We evaluate the row/column-level recognition capability of VLLMs within two specific recognition tasks. Row/column index-based data recognition tasks are designed to assess whether VLLMs can accurately identify row/column elements. Formally, given a table image $I$ with specific cell’s row index $R$ or column index $C$, the VLLMs are required to return the entire row content $[c_{(R,1)},c_{(R,2)},\ldots]$ or the column content $[c_{(1,C)},c_{(2,C)},\ldots]$.
Table-level. We evaluate the overall table-level recognition capability of VLLMs. The visual table size detection recognition task is designed to evaluate whether VLLMs can accurately determine the number of rows and columns in a table, assessing their comprehension of the global structural information.
3.2.2. Baselines
In this study, we evaluate the performance of six VLLMs. For open-source VLLMs, we select Phi-3.5-Vision-Instruct (Phi) and Llama-3.2-90B-Vision-Instruct (Llama) for evaluation. For closed-source VLLMs, we evaluat GPT-4o-mini (GPT-mini), Qwen-VL-Max (Qwen), GPT-4o (GPT) and Gemini-1.5-Pro (Gemini). The specific details of VLLM can be found in the appendix D.
Figure 2. Experimental results of VLLMs for the proposed hierarchical tasks. The tasks evaluated include the following: merged cell detection (MCD), content/index-based cell recognition (CCR, ICR), index-based row/column data recognition (IRDR, ICDR), and visual table size detection (VTSD).
3.3. Benchmark Evaluation Results
Comprehensive evaluations of hierarchical structural recognition tasks are presented in Figure 2. The results show that, among all the VLLMs we selected, GPT and Gemini demonstrate the strongest performance, consistently outperforming the other VLLMs. Furthermore, open-source Llama demonstrates a significant performance gap compared to the closed-source VLLMs. We give some highlights associated with the benchmark results as follows:
Figure 3. Row-column sensitivity analysis of VLLMs on designed hierarchical tasks. The left figure uses Gemini, and the right figure uses GPT.
Row-column Sensitivity Analysis. We found that all models show inconsistent performance between row/column-related tasks. To mitigate the influence of uneven distributions of rows and columns in the original data leading to varying difficulty levels, we further refine the experimental results by selecting samples where the difference between the number of rows and columns does not exceed three. The refined results are presented in Figure 3. The results of row-column sensitivity analysis experiments show that the accuracy for the column-related tasks is higher than for the row-related tasks, suggesting that VLLMs prefer to process column-structured information.
We believe this phenomenon is because columns usually contain more substantial attribute information, allowing VLLMs to locate columns more accurately. Rows tend to have higher similarities, so VLLMs faces more significant challenges locating rows or completing row-related tasks. The differences between rows and columns in visual table size detection further verify this view.
Image Quality Analysis. We conduct in-depth experiments and analysis on the TR task. Detailed experimental results and analysis are presented in Section 5. The results demonstrate that significant bottlenecks remain despite existing VLLMs achieving competitive performance on the TR task. Using simple prompts, the publicly available VLLMs performs relatively well on the SciTSR dataset with higher image quality, scoring only 4.38 TEDS points lower than the leading baseline models. However, when processing the PubTabNet dataset with lower image quality, the performance gap widens significantly, reaching 13.63 TEDS points. This phenomenon indicates that the quality of the input image is a key bottleneck limiting the performance of VLLMs. Section 3.4 provides a more in-depth analysis of this bottleneck.
Table 2. Specific description of the challenges and scenarios in low-quality image inputs.
Challenge Scenario Description
Visual Conditions Blur
The image is out of focus, with details
appearing smeared or indistinct.
Underexposure
The image is too dark, which may
cause the content to be unclear.
Overexposure
The image is overly bright, losing
detail in some regions.
Table Border Quality
Unclear
Borders
The image’s table borders are faint,
blending into the background.
Missing
Borders
The table is without expected
borders or separators.
Thickened
Borders
The table borders are thickened.
Geometric Deformation Tilt 20° The image is tilted at 20° angle.
Tilt 40° The image is tilted at 40° angle.
3.4. Bottleneck Analysis
In this section, we further investigate the performance of VLLMs under varying image quality conditions and assess their visual robustness to these conditions through empirical analysis.
3.4.1. Experimental Setup
To comprehensively evaluate the visual robustness of VLLM, we focus on three distinct visual challenges: the image quality challenge, the table border quality challenge, and the geometric deformation challenge. Specifically, We randomly selected 1,000 samples from the SciTSR (chi2019complicated, ) test set for our experiments and configure our image types according to these challenges. Details of these challenges are provided in Table 2.
Visual Conditions. Visual conditions is a key factor affecting the accuracy of VLLMs in table recognition. To assess it, we systematically analyze the performance of VLLMs under various visual conditions across three scenarios: blur, underexposure, and overexposure. These analyses demonstrate the robustness of VLLMs in handling impaired visual conditions.
Table Border Quality. Table borders indicate structural information and are critical for parsing tasks. We evaluate the impact of border visibility and completeness on table recognition performance by considering scenarios including unclear table borders and missing borders. Additionally, we explore the effect of border changes in the table by thickened table borders.
Geometric Deformation. Geometric deformation caused by viewing angles or operations can disrupt the geometric consistency of tables, which is common in visual table parsing, especially when tables are scanned or photographed. We evaluated the robustness of VLLMs against geometric deformations by testing them under tilt conditions of 20° and excessive tilt conditions of 40°.
3.4.2. Discussion and Analysis
Figure 4 shows the detailed experimental results. We observe that blurring and overexposure can lead to loss of text information in table images and inaccurate fine-grained text recognition, thus negatively impacting the performance of table content recognition. In contrast, underexposure has little effect on table recognition performance. Moreover, VLLM demonstrates limited efficacy when processing skewed tables, resulting in a substantial decrease in accuracy under such conditions.
Although table borders are often considered essential for conveying structured information, modifications such as fading or removing these borders have minimal impact on the performance of VLLMs in TR tasks. This result suggests that VLLMs do not heavily rely on borders in TR. VLLMs only pay slight attention to the structural information the borders provide when thickening. Structural information provided by the borders.
Figure 4. Evaluation results of bottleneck scenarios: abbreviations UE (Underexposure), OE (Overexposure), MB (Missing Borders), UB (Unclear Borders), TB (Thickened Borders).
Figure 5. Illustration of a pipeline for table image preprocessing leveraging a toolkit of lightweight vision models.
4. Neighbor-Guided Toolchain Reasoner
Through our in-depth analysis of VLLMs’ performance on benchmark tasks and bottleneck analysis, we have identified that the quality of the input image emerges as a critical factor that constrains the performance of these VLLMs. Consequently, improving input image quality is essential for enhancing VLLMs’ capability to recognize and interpret structured image data more effectively. To address this bottleneck, we propose the NGTR framework.
4.1. Model Overview
NGTR enhances input image quality by applying various tool combinations tailored to scenarios such as low resolution, overexposure, and noise interference, among others. To select a suitable combination of tools, we employ a similarity-based neighbor retrieval module to obtain samples that resemble each input instance. These samples guide the generation of multiple tool invocation plans. Subsequently, the tool invocation experience learning module executes each toolchain plan and generates the corresponding structured data based on the neighbor samples. It evaluates the effectiveness of different plans to select a relatively better plan. In the final processing stage, we utilize a reflection-driven tool selection module. This module integrates iterative tool invocation and dynamic feedback to refine the processing flow. The optimized image resulting from this process is subsequently input into the VLLMs, which utilize their powerful reasoning capabilities to generate the structured data representation of the table.
Table 3. Specific descriptions of built tools in the toolkit.
Tool Descriptions
Border Enhancement The border enhancement tool improves the legibility of tables and their structures by thickening the border lines in the image. This process enhances the structural information features of tabular data.
Image Upscaling Image upscaling optimize image resolution to improve visual quality. This technique is commonly employed to repair and enhance blurry images.
Noise Reduction The noise reduction tool enhances image quality by adjusting brightness and contrast to mitigate noise interference and underexposure issues.
Binarization The binarization tool converts images to black and white, highlighting key features for easier extraction.
Detection and Cropping This tool identifies table regions within an image and crops them into independent segments. It is particularly suitable for processing images of tables embedded in complex backgrounds.
4.2. Toolkit Preparation
Inspired by the conclusions of the Bottleneck Analysis (Section 3.4), we employ five distinct tools to address various scenarios and potential issues that may arise in the table recognition task. These tools are shown in Table 3. By combining different tools, NGTR effectively addresses various challenging situations. For example, when the table border is faint, the VLLMs can invoke the border enhancement tool to strengthen the structural information by thickening the table border. Similarly, when the table occupies only a portion of the image, the VLLMs can invoke the table detection and cropping tool to identify and crop the relevant table area, thereby reducing noise interference.
4.3. Similarity-based Neighbor Retrieval
Neighbor retrieval methods enable VLLMs to retrieve similar neighbor samples, providing richer contextual information. In the NGTR framework, we hypothesize that images with similar features exhibit similar results after being processed by the same image preprocessing toolchain. Consequently, the processing results of neighbor samples could guide the selection of a potentially optimal toolchain for test samples. We first retrieve images that are similar to the target task images from a sample set (a subset of the training data containing 100 samples). Subsequently, we employ a prompting template to leverage the VLLM’s planning capabilities to generate tool invocation plans. The retrieval process can be formally described as follows:
(1) $\text{Retrieval}(I^{\text{test}},\mathcal{D}^{\prime},f)=\arg\max\left[f\left(I^{\text{test}},I^{i}\right)\right]_{i=1}^{|\mathcal{D}^{\prime}|},$
where $I^{\text{test}}$ represents an image from the test set, $\mathcal{D}^{\prime}$ denotes a subset of the training dataset, and $f$ is the similarity measurement function. In this paper, we combine the ORB (Oriented FAST and Rotated BRIEF) algorithm with the Hamming distance as $f$ to measure the similarity between images. Specifically, the feature vectors are first extracted using the ORB algorithm for the task image and a candidate example image. Subsequently, the Hamming distance between the feature vectors of the two images is calculated as a measure of similarity. Then, we guide the VLLMs to generate multiple tool invocation plans for the image. The generation process can be formally expressed as follows:
(2) $\text{VLLM}(\mathcal{T},\mathcal{N}(I^{\text{test}}))\rightarrow\{p_{1},p_{2},\ldots,p_{n}\},$
where $\mathcal{T}$ represents the description information set of all available image preprocessing tools, including their functions, applicable scenarios, invocation identifiers, and other relevant details; $\mathcal{N}(I^{\text{test}})$ denotes the neighbor image samples of the test sample $I^{\text{test}}$, along with their associated features, retrieved from the training set; and $\{p_{1},p_{2},\ldots,p_{n}\}$ represents the generated candidate set of plans, which are then used to select an appropriate tool invocation plan.
4.4. Tool Invocation Experience Learning
In this module, we follow a sequential workflow to evaluate the multiple tool invocation plans. First, we execute each tool invocation plan generated by the previous module to obtain multiple processed images:
(3) $I_{p_{i}}=f_{p_{i}}(I),\quad i\in\{1,2,\ldots,n\},\vspace{-0.04in}$
where $f_{p_{i}}$ denotes the image preprocessing process corresponding to the tool invocation plan $p_{i}$. Next, we use a simple prompt template to instruct VLLMs to generate an markup sequence.
(4) $\hat{H}_{P_{i}}=\text{VLLM}(I_{P_{i}},P),$
where $P$ represents the prompt to instruct VLLMs generation.
Subsequently, we evaluate the prediction results based on the example labels. The evaluation process employs the tree edit distance-based similarity (TEDS) metrics to quantify the accuracy of the VLLMs output. The specific details of TEDS can be found in the appendix E. By following this process, we calculate a quantitative score for each toolchain, enabling selection of a suitable plan.
4.5. Reflection-driven Tool Selection
Although the tool invocation experience learning module provides a high-quality plan, mindlessly applying the tool invocation plan to new samples may result in the loss of critical information in the image, thereby affecting the accuracy of the final result. To address this, we introduce the reflection-driven tool selection module during the execution phase to refine the processing flow, reduce information loss, and thereby improve recognition accuracy. The formalized expression of the reflection module is as follows:
Let $I^{(t-1)}$ denote the image before the $t$-th operation and $I^{(t)}$ denote the image after the $t$-th operation. The VLLMs computes a decision value $\gamma^{(t)}$ to determine whether to accept the operation:
(5) $\gamma^{(t)}=\text{reflect}(I^{(t-1)},I^{(t)}),$
where $\gamma^{(t)}$ is a binary decision indicating the quality change between the before and after images. The function $\text{reflect}(\cdot)$ evaluates the difference in quality. If $\gamma^{(t)}=1$, the operation is considered successful; otherwise, if $\gamma^{(t)}=0$, the operation is rejected, and the process proceeds to the next step.
This step-by-step module enhances the interaction between the VLLMs and the target image, ensuring the accuracy of the final task outcome. More importantly, introducing this module enables downstream researchers and developers to flexibly customize and expand the toolkit without worrying about the impact of poorly performing expanded tools on the final results, thereby significantly improving the versatility and transferability of our framework. In the last step, we use a simple prompt template to instruct VLLM to generate a markup sequence and obtain the result of table recognition.
5. Experiments
5.1. Experimental Setup
5.1.1. Datasets
In this study, we utilize three widely-used table recognition datasets: SciTSR (chi2019complicated, ), PubTabNet (zhong2020image, ), and WTW (long2021parsing, ), each offering unique characteristics and challenges, as shown in Table 4. SciTSR is a dataset comprising tables extracted from the scientific literature, and the image quality in this dataset is relatively high. In contrast, the image resolution of PubTabNet is 72 pixels per inch, and its overall image quality is relatively low. WTW contains images collected from the wild, introducing a variety of extreme cases, such as tilt, blur, and table curvature, which significantly increase the task’s difficulty. These datasets encompass diverse table types as well as various unique visual challenges, providing a robust foundation for benchmarking in this study.
Table 4. Characteristics of selected table recognition datasets.
Dataset Tables Source
Markup Sequence
Annotations
Textual
Annotations
Year
SciTSR 15K ArXiv $\times$ $\checkmark$ 2019
PubTabNet 568K PubMed $\checkmark$ $\checkmark$ 2020
WTW 14.6K Wild Scenarios $\times$ $\times$ 2021
Table 5. Performance comparison of methods on the SciTSR, PubTabNet, and WTW datasets. ”-” indicates the method’s lack of results (specific reasons are provided in the implementation details). Best scores in the lightweight OCR model category are highlighted in blue, while best scores in the prompt tuning in VLLMs category are highlighted in green.
Dataset Metrics Lightweight OCR Model Prompt Tuning in VLLMs
EDD LGPMA LORE Phi GPT-mini Qwen Llama Gemini GPT-4o
direct NGTR $\Delta$ direct NGTR $\Delta$
SciTSR TEDS - 95.08 - 66.18 87.18 89.40 87.24 90.15 91.07 +0.92 90.70 92.58 +1.88
TEDS-Struct - 96.24 97.22 71.56 92.03 93.06 92.31 93.73 95.09 +1.36 94.20 95.78 +1.58
PubTabNet TEDS 89.67 94.63 - 49.92 58.68 52.53 79.04 81.00 84.80 +3.80 74.46 85.03 +10.57
TEDS-Struct - 96.70 96.94 57.65 73.00 63.90 87.64 85.28 89.30 +4.02 84.91 92.31 +7.40
WTW TEDS-Struct - - 93.86 - 31.72 - 32.87 42.62 44.68 +2.06 40.01 52.03 +12.02
5.1.2. Baselines
We select six Vision Large Language Models (VLLMs), including Phi, Llama, GPT-mini, Qwen, GPT, and Gemini, as baseline models for comparison. These models’ selection and implementation details are described in appendix D. Additionally, we select three representative deep learning-based methods as baselines for comparison: EDD (zhong2020image, ) based on sequence modeling, LGPMA (qiao2021lgpma, ) based on cell bounding box detection, and LORE (xing2023lore, ) based on cell point center detection.
5.1.3. Evaluation Metrics
As for the evaluation metrics of TR, we use a similarity metric based on Tree-Edit Distance (TEDS) (zhong2020image, ) and the structure-aware TEDS-Struct metric. The specific details of TEDS can be found in the appendix E. We employ two evaluation metrics for the hierarchical tasks described in Section 3.2.1: accuracy (ACC) and micro-averaged F1 score (F1-score). Specifically, ACC is used to evaluate cell-level tasks (excluding merged cell detection) and table-level tasks; the F1-score is utilized for row- and column-level tasks and merged cell detection.
5.1.4. Implementation Details
For the PubTabNet dataset, since the test set labels are not public, we choose to use the validation set for evaluation. We randomly select 1,500 images from the validation set of PubTabNet as research samples. The experimental results show that the EDD and LGPMA baseline model’s performance on this sample is consistent with its initial reported results, which shows that our sampling is well-representative. For the SciTSR and WTW datasets, we use their complete test sets for evaluation. Since the WTW dataset does not provide content information for table recognition, we do not report its TEDS scores.
For LORE, since it is mainly aimed at table structure recognition but not table content recognition, we only report its performance scores for table structure recognition. As for EDD, since its model training requires a large amount of end-to-end annotated data, and SciTSR and WTW lack corresponding labeled data, its performance on these datasets has not been evaluated. For LGPMA, since the model relies on table content for training and the WTW dataset does not provide table content labels, we did not evaluate its performance on the WTW dataset.
5.2. Main Results Analysis
Tables 5 show our benchmark results on the table recognition and table structure recognition tasks. Based on the experimental results, we draw the following insights:
5.2.1. Performance Analysis of NGTR
As shown in Table 5, the experimental results compare our NGTR framework with baseline methods. We use GPT and Gemini, which performs well in benchmark tests, as the framework backbone. The main results show that our framework achieves significant performance improvements on the PubTabNet dataset, mainly attributed to our framework’s enhanced VLLMs robustness when dealing with low-quality inputs. On the SciTSR dataset, our framework also outperforms all VLLMs baselines, further verifying our framework’s effectiveness.
In the WTW dataset, our method and all VLLM-based baselines demonstrate suboptimal performance compared to traditional lightweight baseline methods. However, despite these challenges, our approach significantly enhances the effectiveness of VLLMs, suggesting that our framework imparts a degree of robustness to VLLMs in complex environments. Nevertheless, it highlights persistent challenges that remain difficult to address in VLLMs. Section 5.2.4 presents a more detailed analysis of VLLMs performance on the WTW dataset.
5.2.2. Open-source and Proprietary VLLMs
The advanced open-source VLLMs demonstrate capabilities comparable to proprietary VLLMs in this task. As a representative of open-source VLLMs, Llama exhibits outstanding performance, particularly in the table structure evaluation task (TEDS-Struct) on the PubTabNet dataset. Llama achieves a relatively better result, surpassing GPT by 2.73 points and Gemini by 2.36 points. These results confirm the exceptional performance of Llama but also underscore the potential of open-source VLLMs for further research.
5.2.3. Model Scale
We observe that lightweight VLLMs face significant performance limitations in specific tasks. For instance, Phi, a lightweight VLLM with a parameter size of 4.2B, demonstrates consistently low performance across all evaluations. While another lightweight VLLM, GPT-mini, achieves commendable results on the simpler SciTSR dataset, its performance declined significantly on the more challenging PubTabNet dataset, dropping 33% and 21% in TEDS and TEDS-Struct metrics, respectively.
We attribute this phenomenon to two key factors. First, the TR task requires VLLMs to possess fine-grained detection capabilities and the ability to accurately recognize and extract structured information—a difficult challenge. Due to their relatively small parameter sizes and limited computational power, lightweight VLLMs struggle to achieve the required recognition accuracy and input robustness. Second, the TR task demands the generation of accurate markup sequence representations of tables, imposing stringent requirements on the VLLMs’ generation capabilities. We observe that lightweight VLLMs often fail to produce parsable markup sequences after multiple attempts, mainly when tasked with complex scenarios such as large, intricate tables or ambiguous image inputs. This limitation significantly reduces their final performance scores, a phenomenon not observed in larger-scale VLLMs.
5.2.4. WTW Dataset: A Challenge for VLLMs
Table 5 presents the experimental results on the WTW dataset. These results indicate that prompt tuning based VLLMs methods still exhibit significant gaps compared to traditional lightweight OCR methods, highlighting the challenges VLLMs face when processing datasets with wild scenarios. An in-depth analysis of the experimental findings reveals that the WTW dataset primarily emphasizes table structure information. Performance declines notably when processing tables with numerous empty cells, unevenly distributed text, skewed, rotated, or densely packed text. Further analysis of the outputs suggests that VLLMs tend to ignore cells lacking semantic content. While this behavior helps avoid processing irrelevant data, it also limits their ability to effectively capture the structural information of tables with many blank cells.
In conjunction with the analysis of table borders in Section 3.4, we speculate that VLLMs primarily rely on cell content and its logical position in the image when completing the table recognition task, with relatively limited use of table border information. This dependence restricts the performance of VLLMs in recognizing complex table structures to some extent.
Table 6. Ablation study results on key components of the proposed framework. ”EXP” denotes the tool invocation experience learning module, while ”REF” represents the reflection-driven tool selection module.
Method SciTSR PubTabNet
TEDS TEDS-Struct TEDS TEDS-Struct
NGTR 92.56 95.43 85.03 92.31
w/o EXP 90.33 93.68 80.57 88.40
w/o REF 91.53 94.77 82.08 91.85
5.3. Ablation Study w.r.t Key Components
The results of our ablation experiments on the NGTR framework are shown in Table 6. We randomly select 1,000 images from the SciTSR dataset for ablation studies. We ablate two key components: tool invocation experience learning and reflection-driven tool selection module. Note that we do not ablate the toolkit, as it is the foundational module of our framework and cannot be removed. The results highlight the critical contributions of these components to the overall performance of the NGTR framework.
Effectiveness of Tool Invocation Experience Learning. To evaluate the effectiveness of VLLMs in optimizing toolchains using neighbor samples, we performed an ablation study by removing the tool invocation experience learning module. In this experiment, we only prompted VLLMs to generate a tool invocation plan and applied it directly to the test samples. Without the tool invocation experience learning module, the experimental results showed that while the VLLMs was able to generate a valid tool invocation plan, the lack of effective validation of the generated plan led to a significant performance decline. This further validates the critical role of the tool invocation experience learning module in improving the NGTR framework performance.
Effectiveness of Reflection-driven Tool Selection. To evaluate the effectiveness of introducing the reflection-driven tool selection module for stepwise backtracking validation in the final execution stage, we ablate this module. We apply all the tools in the toolchain to the task images in a single pass, using them as input for the generation stage. Experimental results show that, without the reflection module for stepwise backtracking validation, VLLMs cannot effectively supervise the processing procedure, which may lead to the incorporation of unsuitable tools for the current sample, thereby affecting performance. This performance drop is mainly reflected in the reduction of the TEDS indicator, which may be due to the loss of text details caused by the incorrect noise reduction and binarization application.
Figure 6. Impact of tool count (left) and multi-path count (right) on performance.
5.4. Tool Invocation Analysis
To further investigate the NGTR framework’s preferences in tool invocation, we calculated the tool usage rates for each test dataset (i.e., the proportion of samples that invoke a particular tool among all samples where tools are invoked). The results are presented in Table 7. As shown, there are notable differences in the frequency of tool usage rates across different datasets.
On the PubTabNet dataset, the NGTR framework exhibits a marked preference for the image upscaling tool. This preference is consistent with this dataset’s overall lower image quality characteristic, as improving image resolution can better support subsequent table recognition tasks. For the SciTSR and WTW datasets, the border enhancement tool is invoked more frequently. These two datasets primarily feature tables with complete borders, and enhancing border features enables VLLM to recognize the table structure information more accurately. Moreover, the detection and cropping tool has the highest invocation rate on the WTW dataset, which aligns with the dataset’s characteristics in the wild scenarios. In such images, tables often occupy only a portion of the frame or appear displaced, making this tool effective in locating and extracting the target areas. In summary, the experimental results demonstrate that the NGTR framework adaptively selects appropriate image processing tools based on the characteristics of different datasets, thereby fully showcasing its flexible planning capabilities in cross-dataset scenarios.
Table 7. Tool invocation preference analysis. Values show the percentage of samples using each tool, with bold values marking the dataset where each tool is most frequently used.
Dataset
Image
Upscaling
Border
Enhancement
Binarization
Noise
Reduction
Detection and
Cropping
PubTabNet 93.71 76.42 4.40 15.64 64.39
SciTSR 77.97 90.79 3.70 20.64 54.23
WTW 79.81 86.86 6.65 15.46 82.93
5.5. Hyperparameter Sensitivity Analysis
The NGTR framework contains two core parameters: the maximum length of the toolchain execution plan $L$ and the number of plans generated each time $N$. Based on GPT, Experimental results on the SciTSR dataset are presented in Figures 6.
Our experimental results indicate that a moderate toolchain length achieves an adequate balance between complexity and performance, as excessive toolchain length increases combinatorial complexity and limits processing performance, thereby affecting the framework’s ability to generate high-quality solutions. Similarly, generating a moderate number of execution plans effectively balances solution quality and generation efficiency, whereas generating too few or too many plans slightly reduces performance. Therefore, a moderate toolchain length and number of execution plans can balance complexity and performance well, providing valuable guidance for the tool invocation.
6. Conclusion and Limitation
This paper addressed the under-explored challenge of table recognition using VLLMs in a training-free reasoning paradigm. We proposed the NGTR framework, which enhanced input image quality through lightweight models and neighbor-guided tool invocation strategies. Extensive experiments demonstrated that NGTR significantly improved VLLM-based table recognition performance, addressing critical bottlenecks such as low-quality input images. This work not only established a benchmark for table recognition but also highlighted the potential of VLLMs in advancing table understanding, paving the way for future research and applications.
Despite the strengths of our framework, we acknowledge several limitations that warrant further investigation. Firstly, its performance depends on the underlying toolkit. Secondly, when the available set of neighbor candidates does not sufficiently cover a wide range of scenarios, selecting inappropriate neighbors may lead to suboptimal performance. Nevertheless, we believe the NGTR framework demonstrates strong generalizability, serving as a versatile approach for tool invocation for various domains.
Appendix
Appendix A Logical to Markup Sequence Conversion
Since the baseline methods (including LORE and LGPMA) output logical structures representing cell information, they cannot be directly used for the calculation of TEDS metrics. To address this, we use the following pseudocode to convert the logical structure into markup sequence format.
We divide the entire conversion process into two stages. In the first stage, we perform a preliminary preprocessing of the logical location information, associating the logical positions with the corresponding cell content and storing them in a tabular data matrix. The detailed implementation is shown in Algorithm 1.
Algorithm 1 From Logical Location to Tabular Data Matrix
Input: cells = { $C_{1},C_{2}\dots,C_{K}$
Output: table
1:max_row $\leftarrow$ maximum value of ’end_row’ in cells
2:max_col $\leftarrow$ maximum value of ’end_col’ in cells
3:Initialize table as a array with dimensions (max_row, max_col)
4:for cell in cells do
5:    start_row, end_row, start_col, end_col, content $\leftarrow$ cell
6:    rowspan $\leftarrow$ 1 $+$ end_row -- start_row
7:    colspan $\leftarrow$ 1 $+$ end_col -- start_col
8:    for row = start_row to end_row do
9:         for col $\leftarrow$ start_col to end_col do
10:             if row ==absent$=$ = start_row and col ==absent$=$ = start_col then
[…]
12:                 ”colspan”: colspan, ”content”: content }
13:             else
14:                 table[row][col] $\leftarrow$ ”merged”
15:             end if
16:         end for
17:    end for
18:end for
In the second stage, we traverse the tabular data matrix row by row, gradually converting the stored logical information and cell content into a markup sequence, ultimately generating the output. The specific implementation is detailed in Algorithm 2.
Algorithm 2 From Tabular Data Matrix to Markup Sequence
Input: table
Output: markup
1:Initialize markup $\leftarrow$<table>
2:for row in table do
3:    markup += ”<tr>
4:    for cell in row do
5:         if cell == ”merged” then
6:             continue
7:         else
8:             rsp = cell[”rowspan”]
9:             csp = cell[”colspan”]
10:             markup += <td rowspan $=rsp\text{ colspan}=csp\texttt{>}$
11:             markup += cell[”content”] + ”</td>
12:         end if
13:    end for
14:    markup += ”</tr>
15:end for
16:markup += ”</table>
Appendix B Study of Table Size
In this section, we analyze the scale of the tables involved in the benchmark evaluations. We quantify the size of tables in the SciTSR and PubTabNet datasets by counting the number of cells in each table. Furthermore, this section presents experiments designed to evaluate the impact of table size on table recognition performance. GPT is selected as a representative of VLLMs for these experiments, with the distribution details and experimental results illustrated in Figure 7. In small scale tables, the model demonstrates stable performance on both the SciTSR and PubTabNet datasets. However, as the table size increased, the model’s performance experienced a moderate decline. This indicates that larger tables introduce longer context lengths, thereby affecting the VLLMs’ TR task performance.
Figure 7. Table size distribution and its impact on TR performance in SciTSR (left) and PubTabNet (right) Datasets
Specifically, statistical analysis shows that all tables involved in the experiments do not exceed the VLLMs’ context window after being encoded into markup sequence. The largest table occupies only 10K tokens of context when encoded as an markup sequence, while all VLLMs possess context windows of at least 128K tokens. However, it is important to note that generating well formed markup sequence poses certain challenges to the model’s capabilities. Therefore, when tables are huge and have complex structures, some VLLMs with weaker generative abilities may encounter issues generating parseable markup sequence results, even though the 128K context window is not exceeded.
Figure 8. A case study is conducted using an evaluation sample from PubTabNet
Appendix C Case Study
Figure 8 presents a case study that showcases evaluation samples from the PubTabNet dataset and analyzes the table recognition results using the traditional OCR models LORE and VLLM. The input table image lacks clear row borders, making it challenging for traditional OCR models to locate specific cells accurately. However, VLLM effectively comprehends the hierarchical structure of the table, thereby producing correct recognition results. This case highlights the advantages of VLLM in semantically rich table recognition tasks, demonstrating its superior adaptability and robustness compared to traditional OCR models.
Figure 9. A case study is conducted using an evaluation sample from SciTSR
However, relying on the understanding of the table’s hierarchical structure to perform table recognition tasks can also have negative consequences in certain scenarios. As illustrated in Figure 9, we analyzed the results of using VLLM for table recognition on evaluation samples from the SCITSR dataset. The results indicate that the output of VLLM in table recognition significantly deviates from the correct answers, particularly in the placement of the ”Linear” and ”Kernel” cells in the table. Specifically, VLLM tends to misidentify ”Linear” and ”Kernel” as belonging to the first row of the table, treating them as subheaders, even though these cells are located in the table’s upper-left corner. This phenomenon may stem from VLLMs’ excessive reliance on contextual cues, causing the model to be misled by its own reasoning or interpretation mechanisms when encountering atypical table structures, thereby resulting in erroneous hierarchical structural information recognition results.
Appendix D Parameter Settings of VLLMs
We report the parameter settings of six VLLMs. For open-source VLLMs, we select Phi-3.5-Vision-Instruct [^2] https://azure.microsoft.com/en-us/products/phi/ (Phi) and Llama-3.2-90B-Vision-Instruct [^3] https://www.llama.com/ (Llama) for evaluation. For closed-source VLLMs, we evaluate GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ (GPT-mini), Qwen-VL-Max [^5] https://qwenlm.github.io/blog/qwen-vl/ (Qwen), GPT-4o [^6] https://openai.com/index/hello-gpt-4o/ (GPT) and Gemini-1.5-Pro  [^7] https://deepmind.google/technologies/gemini/pro/ (Gemini). Specifically, when generating multiple tool invocation plans, we set the temperature parameter to 0.8 to encourage the generation of more diverse tool invocation plans. For other experiments, the temperature parameter was kept at 0 to ensure the stability of the experimental results. Additionally, the top_p parameter was set to 0.2, and the n_samples parameter was set to 1.
Appendix E Details of TEDS Metric
TEDS is designed to measure the similarity between two tree structures. Specifically, TEDS is used to calculate the similarity between the HTML tree of the real tag and the HTML tree of the predicted tag. When applying the TEDS metric, the HTML table format must first be converted into a tree structure. The similarity is then calculated using the following formula:
(6) $\vspace{-0.04in}\text{TEDS}(T_{a},T_{b})=1-\frac{\text{EditDist}(T_{a},T_{b})}{\max(|T_{a}|,|T_{b}|)},$
where $T_{a}$ and $T_{b}$ represent the tree structures of the tables in HTML format. $\text{EditDist}(T_{a},T_{b})$ denotes the tree-edit distance, and $|T|$ refers to the number of nodes in tree $T$. Additionally, we adopted a modified version of TEDS, referred to as TEDS-Struct. TEDS-Struct is designed to assess the accuracy of table structure recognition without considering the specific results generated by the table content.
Figure 10. Simplified prompt template for VLLMs on table recognition tasks.
Figure 11. Chain-of-Thought prompt template for VLLMs on table recognition tasks.
Appendix F Implementation Details of Benchmark
We present two prompt templates designed to guide VLLM in generating markup sequence outputs for TR tasks in Figures 10 and 11. Prompt 1 is a simplified template that facilitates the execution of TR tasks by VLLM. In contrast, Prompt 2 introduces the concept of Chain-of-Thought by explicitly planning the table recognition process for VLLM. In our experiments, the more complex Prompt 2 achieved a slight improvement in TR tasks performance. However, since our benchmark evaluation aims to reflect VLLM’s recognition capabilities more directly, we employ Prompt 1 as the prompt template in all experiments.
Figures 12-17 illustrate the prompt templates designed for our benchmark evaluation of hierarchical evaluation tasks. Each template corresponds to a specific hierarchical task, as detailed below:
  • Visual Table Size Detection Task. (Figure 12)
  • Merged Cell Detection Task. (Figure 13)
  • Content-based Cell Recognition Task. (Figure 14)
  • Index-based Cell Recognition Task. (Figure 15)
  • Row Index-based Data Recognition Task. (Figure 16)
  • Column Index-based Data Recognition Task. (Figure 17)
Figure 12. Prompt template for VLLMs on visual table size detection task.
Figure 13. Prompt template on merged cell detection task.
Figure 14. Prompt template for VLLMs on content-based cell recognition task.
Figure 15. Prompt template for VLLMs on index-based cell recognition task.
Figure 16. Prompt template for VLLMs on row index-based data recognition task.
Figure 17. Prompt template for VLLMs on column index-based data recognition task.
Figure 18. Prompt template for generating multiple tool invocation plans in the NGTR framework
Figure 19. Prompt template for the reflection-driven tool utilization module in the NGTR framework
Appendix G Implementation Details of NGTR
Figure 18 illustrates the prompt template utilized by our proposed NGTR framework for generating multiple tool invocation plans. By inputting the target image along with its tool descriptions and identifiers, we leverage the planning capabilities of VLLM to produce several tool invocation plans for subsequent modules. This module encompasses two adjustable hyperparameters: the maximum length $L$ of the toolchain execution plan and the number $N$ of plans generated per iteration.
Figure 19 illustrates the prompt template employed by the NGTR framework within the reflection-driven tool utilization module. By inputting both the pre-processed and post-processed images, we could leverage the discriminative capabilities of VLLM to select the image with superior quality. This selection process enables the further optimization of tool invocation strategies, thereby minimizing the loss of image information caused by inappropriate tool usage.